Consider the case of Black Sheep Broadcasting Company: Black Sheep Broadcasting Company (BSBC) just paid a dividend of $2.88What is the expected dividend yield for Black Sheeps stock today? 0 13.14% O 11.80% 09.44% 0 11.35%
Answer #1

Expected return = risk free rate + beta * market risk premium

value of stock = Present value of dividends + Horizontal value

Horizontal value = dividend next year/(Required return - growth rate)

Dividend one year from now

= 2.88 * (1+20%)

= 2.88 * 1.2

= 3.456

= 3.46

Expected return = 5% + 1.8 * 6%

= 15.8%

Horizontal value = 2.88 * 1.2 * 1.04/(0.158-0.04)

= 30.4596610169

= 30.46

Intrinsic value = 3.46/1.158 + 30.4596610169/1.158

= 29.29

Dividend yield = 3.46/29.29

= 11.80%